jacobrios15:

Factor completely. y2 - 12y + 32 a-(y + 4)(y + 8) b-(y - 4)(y - 8) c-(y + 18)(y + 2) d-(y - 18)(y - 2)

Angle:

Here's a quick way to do this one: A. does 4 + 8 = Negative 12? B. does (-4) + (-8) = Negative 12? C. does 18 + 2 = Negative 12? D. does (-18) + (-2) = Negative 12?

jacobrios15:

A.no B.no c.no d.no

Angle:

I think one of them should definitely equal negative 12 what did you get for each?

jacobrios15:

b. equals negative 12

Angle:

tah dahhhh

jacobrios15:

so its b?

Angle:

yup!